OpenStudy (anonymous):

A package contains 16 candy canes, 6 of which are cracked. If 3 are selected, find the probability of getting exactly three cracked candy canes.

OpenStudy (kirbykirby):

\[ \frac{6}{16}\times\frac{5}{15}\times\frac{4}{14}\]